OpenStudy (anonymous):

What is the interquartile range of the following data set? 300, 785, 670, 187, 760, 724 460 598 25 424

OpenStudy (anonymous):

Are 460 598 25 424 part of the data set?

OpenStudy (anonymous):

no those are the answers its multiple choice

OpenStudy (anonymous):

187, 300, 670, 724, 760, 785 Q1 = 300 Q3 = 760 Interquartile range = Q3 - Q1 = 760 - 300 = 460

OpenStudy (amistre64):

it would be useful if you actually participate in the solution process instead of reposting the same question. http://openstudy.com/users/dougy#/updates/51deba09e4b076f7da4140c5